Mel and Ann were having breakfast at a restaurant. Mel paid $3.50 for 3 eggs and 1 sausage patty. Ann paid $5.25 for 2 eggs and 3 sausage patties. Find the cost of each egg and each sausage patty.
Found 2 solutions by rfer, Krystal77:
Answer by rfer(10432)

3e+1s=3.50, s=3.50-3e
2e+3s=5.25
2e+3(3.50-3e)=5.25
2e+10.50-9e=5.25
-7e=-5.25
e=.75
--------------------
3(.75)+s=3.50
2.25+s=3.50
s=1.25
---------------------
proof
2(.75)+3(1.25)=5.25

Answer by Krystal77(1)

Eggs: E
Sasage: S
3.50=3e+s s=3.50-3e
5.25=2e+3s
Substitution:
5.25=2e+3(3.50-3e)
5.25=2e+10.5-9e
0= 5.25-7e
7e=5.25
e= .75
One egg is $0.75

s=3.50-3(.75)
s=3.50-2.25
s= $1.25
One Sasage: 1.25
One Egg: .75